# Normalized mutual information is a biased measure for classification and community detection

### v2

@@ II.2 Normalization of the mutual information
A fundamental difficulty with mutual information as a measure of similarity is that its range of values depends on the particular application, which makes it difficult to say when a value is large or small. Is a mutual information of 10 a large value? Sometimes it is and sometimes it isn’t, depending on the context. To get around this problem one commonly normalizes the mutual information so that it takes a maximum value of 1 when the candidate labeling agrees exactly with the ground truth. There are a number of ways this can be achieved but, as we show here, they are not all equal. In particular, some, including the most popularly used normalization, can result in biased results and should, in our opinion, be avoided. In its place, we propose an alternative, unbiased normalized measure.
@@
The most popular normalized measure, commonly referred to simply as the “normalized mutual information,” uses the plain mutual information $I_{0}(c;g)$ as a base measure and normalizes it thus:
$\text{NMI}_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}$
$=\frac{I_{0}(c;g)+I_{0}(g;c)}{I_{0}(c;c)+I_{0}(g;g)}.$ (24)
@@
This measure has a number of desirable features. Because of the inequalities in (16), its value falls strictly between zero and one. And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript “$(S)$,” for symmetric).
Figure 2: An example of how normalization can impact which labeling is preferred by a mutual information measure. For the ground-truth labeling $g$ (top), the standard unnormalized mutual information $I_{0}(c,g)$ of Eq. (18) gives different scores to the two candidate labelings $c_{A}$ and $c_{B}$, the former receiving a higher score than the latter. By definition, the asymmetrically normalized mutual information will always agree with the unnormalized measure, but the symmetrically normalized measure may not, as is the case here: in this example the symmetric measure scores candidate $c_{B}$ higher. Note that the network itself plays no role here—it is included only as a visual aid.
@@
We can extend the notion of symmetric normalization to any other base measure of mutual information $I_{X}(c;g)$ by writing
@@
All such measures, however, including the standard measure of Eq. (24), share a crucial shortcoming, namely that the normalization depends on the candidate labeling $c$ and hence that the normalized measure can prefer a different candidate labeling to the base measure purely because of the normalization. Figure 2 shows an example of how this can occur. Two candidate labelings $c_{A}$ and $c_{B}$ are considered for the same ground truth $g$. Under the unnormalized mutual information of Eq. (13), candidate A receives a higher score than candidate B, but under the normalized measure of Eq. (24) the reverse is true. This behavior is due to the difference in entropy between the two candidate divisions, the one on the right having larger entropy than the one on the left, which increases its normalization factor and correspondingly decreases the normalized mutual information.
We argue that the unnormalized measure is more correct on this question, having a direct justification in terms of information theory. The purpose of the normalization is purely to map the values of the measure onto a convenient numerical interval, and should not change outcomes as it does here. Moreover, different symmetric normalizations can produce different results. For instance, if one normalizes by $\max(I_{0}(c;c),I_{0}(g;g))$ in Fig. 2 then candidate $c_{A}$ is favored in all cases.
These issues are unavoidable when using a symmetric normalization scheme. In any such scheme the normalization must depend on both $c$ and $g$ and hence can vary with the candidate labeling. However, if we drop the requirement of symmetry then we can normalize in a straightforward way that avoids these issues. We define the asymmetric normalization of any base measure $I_{X}$ as
@@
This definition still gives $I_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$. In fact, Eq. (29) is the only way to normalize such that $I_{X}^{(A)}(g;g)=1$ while simultaneously ensuring that the preferred candidate is always the same as for the base measure. Thus this measure also removes any ambiguity about how one should perform the normalization. Loosely, this asymmetrically normalized mutual information measures “how much $c$ tells us about $g$ as a fraction of all there is to know about $g$.”
The amount of bias inherent in the symmetrically normalized measure when compared with asymmetric one can be quantified by the ratio between the two:
@@
Note how this expression results in a particular bias against complex candidate labelings $c$ that have higher entropy $H(c)$ than the ground truth. Traditional mutual information measures are particularly problematic when the candidate is a refinement of the ground truth, meaning the candidate groups are subsets of the ground-truth groups. The example given in Section II.1, in which the candidate labeling puts every node in a group on its own, is an extreme instance of this. Equation (31) shows that, to some extent, the traditional symmetric normalization corrects this issue: a candidate $c$ that takes the form of a refinement of $g$ will have $H(c)>H(g)$ and hence the normalized mutual information will be penalized.
@@
We argue, however, that this is not the best way to address the problem and that the correct approach is instead to use the reduced mutual information. The reduced mutual information also penalizes such cases, but it does so in a more principled manner that directly addresses the root cause of the problem, rather than merely penalizing complex candidate labelings in an ad hoc manner as a side-effect of normalization. We will see examples of this in Section III, where we apply both measures to community detection in networks and find that, although the symmetrically normalized traditional measure is biased, it is in some cases fortuitously biased in the right direction, although it is still problematic in some others.
An obvious downside of asymmetric normalization is the loss of the symmetry in the final measure. In the most common applications of normalized mutual information, where labelings are evaluated against a ground truth, an inherently asymmetric situation, the asymmetric measure makes perfect sense, but in other cases the lack of symmetry can be undesirable. Embedding and visualization methods that employ mutual information as a similarity measure, for example, normally demand symmetry [14]. And in cases where one is comparing two candidate labelings directly to one another, rather than to a separate ground truth, then a symmetric measure may be preferable.
@@
Even in this case, however, the asymmetric measure may sometimes be the better choice, as discussed in the introduction. For instance, when one labeling $c_{1}$ is a refinement of the other $c_{2}$, the information content is inherently asymmetric: $c_{1}$ says more about $c_{2}$ than $c_{2}$ does about $c_{1}$. An explicit example of this type of asymmetry is shown in Figure 3, where we consider two labelings of 27 objects. The left labeling $c_{1}$ is a detailed partition of the objects into nine small groups while the right labeling $c_{2}$ is a coarser partition into only three groups, each of which is an amalgamation of three of the smaller groups in $c_{1}$. Because of this nested relationship, it is relatively easy to transmit $c_{2}$ given knowledge of $c_{1}$ but more difficult to do the reverse. This imbalance is reflected in the asymmetrically normalized mutual information values in each direction (top and bottom arrows in the figure), but absent from the symmetric version (middle row).
@@
where the mutual information $I(c;g)$ is quantified as in Eq. (12). This measure correctly accounts for the information contained in the contingency table, returns 0 in expectation when $c$ is uncorrelated with the ground truth and 1 when $c=g$, and always favors the same labeling as the unnormalized measure.
@@
Figure 3: Values of the various normalized mutual information measures we consider, between two example partitions of the same set of 27 objects. The left partition $c_{1}$ divides the objects into nine groups of three objects each and is a refinement of the right partition $c_{2}$, which divides them into only three groups, aggregating those on the left. The arrows indicate the direction of the comparison and the accompanying notations give the values of the mutual information measures. All three asymmetrically normalized measures capture the intuition that the partition $c_{1}$ tells us more about $c_{2}$ than $c_{2}$ tells us about $c_{1}$.
III Example application:
Community detection
@@ III Example application:
As an example of the performance of the various measures discussed above, we describe in this section the results of an extensive series of tests in which the measures are used to score the output of algorithms for network community detection. In these tests we use the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model [15] to generate random test networks with known community structure and realistic distributions of node degrees and group sizes, and then attempt to recover that structure using a variety of standard community detection algorithms, quantifying the accuracy of the recovery with six different measures: symmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information, and asymmetrically normalized versions of the same. A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks [10, 11, 12, 13], but using only the symmetrically normalized, non-reduced mutual information as a similarity measure. Our results indicate that this measure can produce biased outcomes and we recommend the asymmetric reduced mutual information instead.
@@
The LFR model contains a number of free parameters that control the size of the networks generated, their degree distribution, the distribution of community sizes, and the relative probability of within- and between-group edges. (We give further details on the LFR generative process in Appendix A.) We find that the distributions of degrees and community sizes do not significantly impact the relative performance of the various algorithms tested and that performance differences are driven primarily by the size $n$ of the networks and the mixing parameter $\mu$ that controls the ratio of within/between-group connections, so our tests focus on performance as a function of these variables.
@@
Figure 4: The colors in each panel in the center of this figure indicate which of three community detection algorithms (InfoMap (red) and modularity maximization with $\gamma=1$ (green) and $\gamma=10$ (blue)—see key on right) are best able to find the known communities in a large set of LFR benchmark networks, according to the six mutual information measures we consider. Mixtures of red, green, and blue denote the proportions of test cases in which each algorithm performs best. Regions in gray indicate parameter values for which no algorithm achieved a positive mutual information score.

### v3

@@ 2.1 Normalization of the mutual information
A fundamental difficulty with mutual information as a measure of similarity is that its range of values depends on the particular application, which makes it difficult to say when a value is large or small. Is a mutual information of 10 a large value? Sometimes it is and sometimes it isn’t, depending on the context. To get around this obstacle one commonly normalizes the mutual information so that it takes a maximum value of 1 when the candidate labeling agrees exactly with the ground truth. There are a number of ways this can be achieved and, as we show here, they are not all equal. In particular, some, including the most popularly used normalization, can result in biased results and should, in our opinion, be avoided. In its place, we propose an alternative, unbiased normalized measure.
@@
The most popular normalized measure, commonly referred to simply as the normalized mutual information, uses the plain mutual information $I_{0}(c;g)$ as a base measure and normalizes it thus:
$\text{NMI}_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}=\frac{I_{0}(c;g)+I_{0}(g;c)}{I_{0}(c;c)+I_{0}(g;g)}.$ (7)
@@
This measure has a number of desirable features. Because of the inequalities in (2), its value falls strictly between zero and one. And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript $(S)$, for symmetric).
Figure 1: Normalization can impact which labeling is preferred by a mutual information measure. In this example 64 objects are split into four equally sized true groups $g$, denoted by their color. Against this ground truth, we compare two candidate labelings, $c_{1}$ and $c_{2}$. The standard unnormalized mutual information $I_{0}(c;g)$ of Eq. (5) reports that labeling $c_{1}$ shares more bits of information with the ground truth than does $c_{2}$. By definition, the asymmetrically normalized mutual information $\text{NMI}^{(A)}$ will always agree with the unnormalized measure, as in this example. The symmetrically normalized $\text{NMI}^{(S)}$, on the other hand, is biased in favor of simpler labelings, which makes it prefer labeling $c_{2}$ in this case.
@@
We can extend the notion of symmetric normalization to any other base measure of mutual information $I_{X}(c;g)$ satisfying the inequality Eq. (2), such as adjusted or reduced mutual information [7], by writing
@@
All such measures, however, including the standard measure of Eq. (7), share a crucial shortcoming, that the normalization depends on the candidate labeling $c$ and hence that the normalized measure can prefer a different candidate labeling to the base measure purely because of the normalization.
Figure 1 shows an example of how this can occur. In this example 64 objects are split into four equally sized groups in the ground-truth labeling $g$, against which we compare two candidate labelings, $c_{1}$ and $c_{2}$. Neither candidate is perfect. Labeling $c_{1}$ identifies the correct number of groups but misplaces certain objects, while labeling $c_{2}$ contains only two groups, each an aggregation of a pair of true groups. Under the unnormalized mutual information of Eq. (5), $c_{1}$ receives a higher score than $c_{2}$, but under the normalized measure of Eq. (7) the reverse is true. This behavior is due to the difference in entropy $I_{0}(c;c)$ between the two candidate divisions. Since the candidate entropy appears in the denominator of the symmetric normalization, Eq. (7), the symmetric measure favors simple labelings with smaller entropy. In this example, candidate $c_{2}$ has only two groups, which results in a smaller entropy and a bias in its favor.
We argue that the unnormalized measure is more correct on this question, having a direct justification in terms of information theory. The purpose of the normalization is merely to map the values of the measure onto a convenient numerical interval, and should not change the outcome as it does here. Moreover, different symmetric normalizations can produce different results. For instance, if one normalizes by $\max(I_{0}(c;c),I_{0}(g;g))$ in Fig. 1 then candidate $c_{1}$ is favored in all cases.
The choice of normalization can also change conclusions even when the number of groups does not vary. Consider the example shown in Fig. 2, in which the ground truth $g$ splits the objects into two equally sized groups. Candidate $c_{1}$ also has two groups of equal size, although each is slightly polluted with the objects of the opposite group. Candidate $c_{2}$ has two groups of different sizes, a small group with one color only and a larger group with objects of both colors. Although both labelings have the same number of groups, candidate $c_{2}$ has a smaller entropy $I_{0}(c;c)$, since it is easier to communicate group identity within $c_{2}$—most objects belong to the larger group. As a result, the symmetrically normalized mutual information prefers $c_{2}$, while the unnormalized measure prefers $c_{1}$.
Figure 2: Another example where the choice of normalization affects which labeling is preferred, this one with the same number of groups in both divisions. Although both candidates $c_{1}$ and $c_{2}$ correctly split the objects into two groups, candidate labeling $c_{1}$ scores higher with respect to both the unnormalized and asymmetrically normalized mutual informations while the symmetric normalization prefers $c_{2}$.
These issues are unavoidable when using a symmetric normalization scheme. In any such scheme the normalization must depend on both $c$ and $g$ and hence can vary with the candidate labeling. However, if we drop the requirement of symmetry then we can normalize in a way that avoids these issues. We define the asymmetric normalization of any base measure $I_{X}$ as
@@
This definition still gives $\text{NMI}_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$. In fact, Eq. (10) is the only way to normalize such that $I_{X}^{(A)}(g;g)=1$ while simultaneously ensuring that the preferred candidate is always the same as for the base measure. Thus this measure also removes any ambiguity about how one should perform the normalization. Loosely, this asymmetrically normalized mutual information measures how much $c$ tells us about $g$ as a fraction of all there is to know about $g$. Asymmetrically normalized measures have appeared in other contexts under the name “coefficient of constraint” or “coefficient of uncertainty” [15, 16, 17], although symmetric normalizations are almost universally used in the model validation context, where they introduce the bias we discuss here.
The amount of bias inherent in the symmetrically normalized measure when compared with the asymmetric one can be quantified by the ratio between the two:
@@
Values of this quantity below (above) 1 indicate that the symmetric measure is biased low (high). Thus, for instance, complex candidate labelings $c$ that have higher entropy than the ground truth will result in a symmetric measure whose values are too low. As discussed in the introduction, traditional mutual information measures are particularly problematic when the candidate is a refinement of the ground truth, meaning the candidate groups are subsets of the ground-truth groups. In that case the traditional measure returns values that are too high. Equation (12) implies that, to some extent, the symmetric normalization will correct this issue: a candidate $c$ that takes the form of a refinement of $g$ will have $H(c)>H(g)$, which will lower the value of the symmetrically normalized mutual information. This may in part explain why these biases have been overlooked in the past: two wrongs have conveniently canceled to (approximately) make a right.
@@
We argue, however, that this is not the best way to address the problem and that the correct approach is instead to use the reduced mutual information. The reduced mutual information also corrects for the case where one labeling is a refinement of the other, but does so in a more principled manner that directly addresses the root cause of the problem, rather than merely penalizing complex candidate labelings in an ad hoc manner as a side-effect of normalization.
We will see examples of these effects in Section 2.2, where we apply the various measures to community detection in networks and find that indeed the traditional symmetrically normalized mutual information is biased. In some cases it is fortuitously biased in the right direction, although it is still problematic in some others.
@@
An obvious downside of asymmetric normalization is the loss of the symmetry in the final measure. In the most common applications of normalized mutual information, where labelings are evaluated against a ground truth, an inherently asymmetric situation, the asymmetric measure makes sense, but in other cases the lack of symmetry can be undesirable. Embedding and visualization methods that employ mutual information as a similarity measure, for example, normally demand symmetry [18]. And in cases where one is comparing two candidate labelings directly to one another, rather than to a separate ground truth, a symmetric measure may be preferable. Even in this latter case, however, the asymmetric measure may sometimes be the better choice, as discussed in the introduction. For instance, when one labeling $c_{1}$ is a refinement of the other $c_{2}$, the information content is inherently asymmetric: $c_{1}$ says more about $c_{2}$ than $c_{2}$ does about $c_{1}$. An explicit example of this type of asymmetry is shown in Figure 3, where we consider two labelings of 27 objects. The left labeling $c_{1}$ is a detailed partition of the objects into nine small groups while the right labeling $c_{2}$ is a coarser partition into only three groups, each of which is an amalgamation of three of the smaller groups in $c_{1}$. Because of this nested relationship, it is relatively easy to transmit $c_{2}$ given knowledge of $c_{1}$ but more difficult to do the reverse. This imbalance is reflected in the asymmetric normalized mutual information values in each direction (top and bottom arrows in the figure), but absent from the symmetric version (middle row).
@@
where the mutual information $I(c;g)$ is quantified as in Eq. (4). This measure correctly accounts for the information contained in the contingency table, returns a negative value when $c$ is unhelpful for recovering the ground truth, returns 1 if and only if $c=g$, and always favors the same labeling as the unnormalized measure. The traditional normalized mutual information possesses none of these desirable qualities.
@@
Figure 3: Values of the various normalized mutual information measures between two example partitions of the same set of 27 objects. The left partition $c_{1}$ divides the objects into nine groups of three objects each and is a refinement of the right partition $c_{2}$, which divides them into only three groups. The arrows indicate the direction of the comparison and the accompanying notations give the values of the mutual information measures. All three asymmetrically normalized measures capture the intuition that the partition $c_{1}$ tells us more about $c_{2}$ than $c_{2}$ tells us about $c_{1}$.
2.2 Copurchasing network of books
As an example of the performance of the various measures discussed above, in this section we use them to score the output of algorithms for network community detection. We consider a selection of networks with known true groups and attempt to recover those groups using various standard community detection algorithms, quantifying the accuracy of recovery with six different measures: the symmetrically and asymmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information.
As a first example, Fig. 4 shows a network compiled by V. Krebs of 105 books about US politics published around the time of the 2004 US presidential election. The true groups $g$ in this case represent the political lean of each book’s content: liberal, conservative, or neutral [19], as determined by Krebs. Network edges connect pairs of books frequently purchased together on Amazon.com, and such books typically are of the same political bent, either both liberal or both conservative, so that the political positions form communities of connected network nodes. Community detection algorithms, which aim to find such communities, should therefore be able to recover political positions, at least approximately, from the network of connections alone. The figure shows the results of applying two of the most popular such algorithms, InfoMap and modularity maximization (with resolution parameter $\gamma=2$), but we find that these algorithms identify quite different partitions of the network, with InfoMap finding six groups where modularity finds nine. We use the normalized mutual information in its various forms to assess how closely each division matches the ground truth.
We find that both the base mutual information measure $I_{X}$ and its normalization considerably impact our assessments. For instance, the asymmetrically normalized traditional mutual information gives a score of $\text{NMI}_{0}^{(A)}(c_{2};g)=0.728$, out of a maximum of 1, for the partition found by modularity maximization, which is over twice the score of $\text{NMI}^{(S)}(c_{2};g)=0.294$ given to the same partition by the symmetrically normalized reduced measure.
The choice of mutual information measure can also change which algorithm is preferred. For the traditional unreduced base measure $I_{0}$, the asymmetrically normalized mutual information favors the groups found by modularity while the symmetric version favors the simpler partition found by InfoMap. The same pattern holds if the adjusted $I_{A}$ is used. The reduced measure $I$, however, prefers InfoMap regardless of normalization, although by a narrower margin in the asymmetric case. As discussed in the previous section, we favor the asymmetrically normalized reduced measure on formal grounds, and view the (often large) deviations from its assessment as unnecessary biases.
Figure 4: Performance of the InfoMap and $\gamma=2$ modularity maximization algorithms in recovering the true groups in the copurchasing network of political books [19]. True and recovered groups are indicated by node color, and similarity between the two is assessed using the mutual information in its traditional $I_{0}$, adjusted $I_{A}$, and reduced II variants, normalized asymmetrically $\text{NMI}^{(A)}$ and symmetrically $\text{NMI}^{(S)}$. The choices of both the base measure and the normalization impact which algorithm is preferred.
2.3 LFR benchmark networks
@@
In one sense the previous example is unusual: there are relatively few cases like this of networks with known ground-truth communities to compare against. To get a more comprehensive picture of algorithm performance, therefore, researchers have turned to synthetic benchmark networks like those generated by the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model [20], which creates networks with known community structure and realistic distributions of node degrees and group sizes. A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks [11, 12, 13, 14], but using only the symmetrically normalized, non-reduced mutual information as a similarity measure. Our results indicate that this measure can produce biased outcomes and we recommend the asymmetric reduced mutual information instead.
@@
The LFR model contains a number of free parameters that control the size of the networks generated, their degree distribution, the distribution of community sizes, and the relative probability of within- and between-group edges. (See Section A in the Supplementary Materials for details of the LFR generative process.) We find that the distributions of degrees and community sizes do not significantly impact the relative performance of the various algorithms tested and that performance differences are driven primarily by the size $n$ of the networks and the mixing parameter $\mu$ that controls the ratio of connections within and between groups, so our tests focus on performance as a function of these parameters.
@@
Figure 5: Performance of three community detection algorithms: InfoMap (red) and modularity maximization with $\gamma=1$ (green) and $\gamma=10$ (blue)—see key on right. The colors in each panel of this figure indicate which of the three is best able to find the known communities in a large set of LFR benchmark networks, according to the six mutual information measures we consider. Mixtures of red, green, and blue denote the proportions of test cases in which each algorithm performs best. Regions in gray indicate parameter values for which no algorithm achieved a positive mutual information score.
2.4 Comparison between variants of the mutual information
Figure 5 summarizes the relative performance of the various mutual information measures in our tests. In this set of tests we limit ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measure which of the three returns the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$. Each point in each of the six panels is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to each of the six measures and, as we can see, the results vary significantly among measures. An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
One consistent feature of all six mutual information measures is the large red area in each panel of Fig. 5, which represents the region in which the InfoMap algorithm performs best. Regardless of the measure used, InfoMap is the best performer on networks with low mixing parameter (i.e., strong community structure) and relatively large network size. For higher mixing (weaker structure) or smaller network sizes, modularity maximization does better. Which version of modularity is best, however, depends strongly on the mutual information measure. The traditional symmetrically normalized mutual information (top left panel) mostly favors the version with a high resolution parameter of $\gamma=10$ (blue), but the asymmetric reduced measure for which we advocate (bottom right) favors the version with $\gamma=1$ (green). (The regions colored gray in the figure are those in which no algorithm receives a positive mutual information score and hence all algorithms can be interpreted as failing.)
These results raise significant doubts about the traditional measure. Consider the lower right corner of each plot in Fig. 5, which is the regime of small network sizes $n$ and high mixing $\mu$ so that the community signal is weak and the noise is high. Here, the adjusted and reduced mutual informations indicate that all algorithms are failing. This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities [21, 22, 23]. The standard normalized mutual information, on the other hand, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization. This occurs because the $\gamma=10$ algorithm finds many small communities and, as discussed in Section 2, a labeling with many communities, even completely random ones, is accorded a high score by a non-reduced mutual information. The presence of this effect is demonstrated explicitly in the Supplementary Materials, Section A.1 for the case of modularity maximization with $\gamma=10$, and this offers a clear reason to avoid the standard measure.
Figure 6: Performance of each of the six community detection algorithms considered here in identifying the known communities in a large set of LFR benchmark networks, as quantified by the asymmetrically normalized reduced information measure $I^{(A)}(c;g)$ proposed in this paper. Gray areas indicate parameter values for which $\text{NMI}<0$, so that the conditional encoding is less efficient than the direct encoding.
The bottom left panel in Fig. 5 shows results for the asymmetrically normalized version of the traditional mutual information, which gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version. This behavior arises for the reasons discussed in Section 2.1—the bias inherent in the symmetric normalization fortuitously acts to partially correct the errors introduced by neglecting the information content of the contingency table. The asymmetric normalization eliminates this correction and hence performs more poorly. The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways as we have seen, but rather to adopt a reduced mutual information measure.
Finally, comparing the middle and right-hand columns of Fig. 5, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences. In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure. This occurs because, as discussed in Section 2 and Ref. [6], the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails—in this regime the candidate and ground-truth labelings are uncorrelated which results in a uniform contingency table. This provides further evidence in favor of using a reduced mutual information measure.
2.5 Comparison between community detection algorithms
Figure 6 shows results for all six algorithms listed in the Supplementary Materials, Section A.2. Examining the figure, we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model. Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better. As noted by previous authors [24], InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime. In this regime one must use another algorithm. Either the modularity or inference method are reasonable options, but modularity has a slight edge, except in a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore. (We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Section A of the Supplementary Materials.)
Thus—always assuming the LFR benchmark is a good test of performance—our recommendations for the best community detection algorithm are relatively straightforward. If we are in a regime where InfoMap succeeds, meaning it finds more than one community, then one should use InfoMap. If not, one should use standard modularity maximization with $\gamma=1$. That still leaves open the question of how the modularity should be maximized. In our studies we find the best results with simulated annealing, but simulated annealing is computationally expensive. In regimes where it is not feasible, we recommend using the Leiden algorithm instead. (Tests using other computationally efficient maximization schemes, such as the Louvain and spectral algorithms, generally performed less well than the Leiden algorithm.)
3 Discussion
In this paper we have examined the performance of a range of mutual information measures for comparing labelings of objects in classification, clustering, or community detection applications. We argue that the commonly used normalized mutual information is biased in two ways: (1) because it ignores the information content of the contingency table, which can be large, and (2) because the symmetric normalization it employs introduces spurious dependence on the labeling. We argue in favor of a different measure, an asymmetrically normalized version of the reduced mutual information, which rectifies both of these shortcomings.
4 Data availability
All datasets analyzed here have been deposited in the Zenodo database under accession code 10.5281/zenodo.17211478, except for the synthetic LFR benchmark networks, which were randomly generated using the open-source NetworkX package as described in the Supplementary Materials, Section A.
Code availability
Python code implementing the mutual information measures discussed in this paper is available at https://github.com/maxjerdee/clustering-mi or [25].
Acknowledgments
The authors thank Samin Aref for useful comments and feedback. This work was supported in part by the US National Science Foundation under grants DMS–2005899 and DMS–2404617 (MN), the National Science Foundation of China through Young Scientist Fund Project grant 12405044 (AK), and computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
Author contributions
All authors (MJ, AK, and MN) conceptualized the research and wrote and edited the paper. Code development and data analysis were performed by MJ. All authors have read the final manuscript.
Competing interests
The authors declare no competing interests.
Appendix A Benchmarking
In this appendix we give some additional details of our community detection tests.
A.1 Numbers of groups
As discussed in Section 2.4, the traditional (non-reduced) mutual information favors the $\gamma=10$ generalized modularity maximization algorithm over other algorithms across many of the tests reported in Fig. 5. However, a closer inspection of the partitions found by this algorithm reveals that it drastically overestimates the number of communities. The number of groups inferred by each algorithm for each value of the mixing parameter $\mu$ is shown for $n=3200$ in Fig. 7. The traditional mutual information has a bias towards labelings with an excessive number of groups, which causes it to prefer the $\gamma=10$ algorithm in this regime, while the reduced and adjusted mutual information measures prefer simple modularity maximization ($\gamma=1$).
Figure 7: The number of groups inferred by each of the six algorithms in Fig. 5 for LFR benchmark networks with $n=3200$ nodes and a range of values of the mixing parameter $\mu$. The ground-truth number of groups is shown in black. The InfoMap algorithm (red) generates an accurate number of groups for values of $\mu$ up to about 0.5, but beyond this point it erroneously places all nodes in a single group. Standard modularity maximization with resolution parameter $\gamma=1$ (green) underestimates the number of groups, perhaps because of the resolution limit on the detection of small groups, but not as severely as the number is overestimated when $\gamma=10$ (blue).
A.2 Community detection algorithms